# SHyPar: A Spectral Coarsening Approach to Hypergraph Partitioning

### v2

@@ III-B2 Local clustering algorithms
The proposed flow-based clustering algorithm in [67] is strongly local by expanding the network around the seed nodes $\mathcal{C}$, which obviously benefits the coarsening framework: (1) applying the max $s$-$t$ flow, min $s$-$t$ cut problem on the local neighborhood of the seed nodes restricts node-aggregation locally and keeps the global hypergraph structure intact; (2) such a local clustering scheme significantly improves the algorithm efficiency due to the small-scale input dataset.
@@ III-B3 Flow-based Local Clustering in HyperSF
First, we apply HyperEF (Algorithm 2) to the hypergraph $H=(V,E,w)$ to compute a coarsened hypergraph $H^{\prime}=(V^{\prime},E^{\prime},w^{\prime})$ and identify isolated nodes, denoted by $\mathcal{C}$. Leveraging the introduced flow-based clustering in [67], HyperSF constructs a sub-hypergraph $H^{\prime}_{L}$ by iteratively expanding the hypergraph around the seed node set $\mathcal{C}$ and then repeatedly solve the hypergraph cut problem to minimize HLC until no significant changes in local conductance are observed. Define $E^{\prime}(S)=\cup_{v^{\prime}\in V^{\prime},v^{\prime}\in\{e^{\prime}\}_{e^{\prime}\in E^{\prime}}}E^{\prime}(v)$ for any set $S\subseteq V^{\prime}$ and let $H^{\prime}_{\varsigma}=(V^{\prime}\cup\{s,t\},E^{\prime}\cup E^{\prime st})$, where $E^{\prime st}$ denotes the terminal edge set. HyperSF aims to construct a sub-hypergraph $H^{\prime}_{L}$ to replace $H^{\prime}_{\varsigma}$ that minimizes HLC by repeatedly solving a local version of (21). To this end, HyperSF sets up an oracle to discover a set of best neighborhood vertices for a given vertex $v^{\prime}$:
@@ III-C Algorithm Complexity for Spectral Coarsening
III-D Hypergraph Partitioning with Spectral Coarsening
III-D1 Multilevel spectral coarsening
The spectral hypergraph coarsening method constructs a sequence of successively smaller (coarser) but spectrally-similar hypergraphs via repeatedly contracting groups of vertices (clusters). Initially, the node spectral embeddings are computed by applying the Krylov subspace method in linear time to obtain the vectors of various spectra. The optimization-based effective-resistance estimation formulation is extended to the hypergraph settings by leveraging the nonlinear quadratic operator of the hypergraphs. Then, a low-resistance-diameter hypergraph decomposition method determines the highly connected vertices in multi-dimensional spectral space.
We introduce a new rating function based on the effective resistance formula, which selects a pair of vertex ($p,q$) from hyperedges that have a large number of heavy nets with low effective resistance.
$r(p,q)=\sum_{e=(p,q)\in E}{\frac{w(e)}{|R_{e}|-1}},$ (23)
where $r(p,q)$ is the rating of the vertex pair, $w(e)$ is the weight of hyperedge $e$, and $|R_{e}|$ is the hyperedge effective resistance calculated by Eq. (17).
III-D2 Multilevel coarsening with HyperEF
The resistance-based hypergraph clustering algorithm allows transferring hypergraph structural information through the levels by assigning a weight (that is equal to the effective-resistance diameter of each cluster evaluated at the previous level) to the new corresponding vertices. As a result, the effective hyperedge resistance at a coarse level depends on not only the evaluated effective resistance computed at the current level but also the results transferred from all the previous (finer) levels.
Figure 8: Multilevel spectral coarsening.
As an example, Figure 8 illustrates the two-level coarsening of a hypergraph with 9 vertices and 4 hyperedges. Initially, the vertices are clustered into 4 clusters $C_{1}$, $C_{2}$, $C_{3}$, $C_{4}$, and the vertices within each cluster are aggregated to create a coarser hypergraph with 4 vertices and 3 hyperedges. In the next level, 4 vertices are grouped into 2 clusters $C^{\prime}_{1}$ and $C^{\prime}_{2}$ that form a coarser hypergraph with 2 vertices and 1 hyperedge.
III-D3 Hypergraph clustering with HyperSF
The flow-based clustering algorithm introduced in [67] is used as a refinement technique to enhance the computed clusters by HyperEF. Sorting the clusters by resistance diameter, HyperSF begins with those that have the lowest values. It then applies a flow-based clustering technique to form new clusters by merging nearby clusters with low resistance diameters within the same neighborhood. It is important to note that clusters with low resistance diameters typically indicate a strong connection among the nodes within the cluster, and therefore, merging these nodes is unlikely to significantly alter the global structure of the hypergraph. Ultimately, HyperSF aims to produce clusters characterized by low hypergraph local conductance (HLC) to preserve the spectral properties of the original hypergraph.
In our approach, we perform community detection on hypergraphs by translating this challenge into a hypergraph clustering problem. Our algorithm identifies the node cluster with the smallest effective-resistance diameter and refines it by applying a strongly-local flow-based method (HyperSF). This is followed by a community-aware coarsening phase that applies the hypergraph coarsening algorithm to each community separately.
@@ IV Experimental Validation
There are many applications related to hypergraph partitioning. This paper focuses on comprehensively evaluating the performance of the proposed hypergraph partitioning framework for increasingly important applications related to integrated circuits computer-aided design. Both the solution quality and runtime efficiency will be carefully assessed by testing them on a wide range of public-domain data sets. The developed open-source software packages will also be made available for public assessment.
@@ IV-B Hypergraph Partitioning with Spectral Coarsening
We compared SHyPar with leading hypergraph partitioners hMETIS [2], SpecPart [43], KaHyPar [45], and MedPart [69] using two sets of publicly available benchmarks: the ISPD98 VLSI Circuit Benchmark Suite [68] and the Titan23 Suite [70]. The details of these benchmarks are outlined in Table III and Table IV. All tests were conducted on a server equipped with Intel(R) Xeon(R) Gold 6244 processors and an NVIDIA Tesla V100S GPU with 1546GB of memory.
@@ IV-B2 SHyPar Performance on ISPD98 Benchmarks
TABLE III: Statistics of ISPD98 VLSI circuit benchmark suite. The best results among all the methods are colored red.
@@ IV-B3 SHyPar Performance on Titan23 Benchmarks
Building on the findings from the ISPD98 benchmarks, our research extended to the Titan23 benchmarks, notable for their high-degree hyperedges. Table IV details the results on the Titan23 benchmarks, where SHyPar dramatically outperforms hMETIS, achieving a 12% improvement for $\epsilon$ = $2\%$ and an impressive 25% for $\epsilon$ = $20\%$. Also SHyPar outperforms SpecPart, achieving a 3% improvement for $\epsilon$ = $2\%$ and 1% for $\epsilon$ = $20\%$. In cases such as sparcT2_core, SHyPar even exceeds the best published results by up to 15%; these notable achievements are highlighted by underlining.
TABLE IV: Statistics of Titan23 benchmark suite and cut sizes of different approaches. The best results among all the methods are colored red.

### v3

@@ III-B2 Local clustering algorithms
The proposed flow-based clustering algorithm in [54] is strongly local by expanding the network around the seed nodes $\mathcal{C}$, which benefits the coarsening framework in two ways: (1) applying the max $s$-$t$ flow, min $s$-$t$ cut problem on the local neighborhood of the seed nodes restricts node-aggregation locally and keeps the global hypergraph structure intact; (2) such a local clustering scheme significantly improves the algorithm efficiency due to the small-scale input dataset.
@@ III-B3 Flow-based Local Clustering in HyperSF
First, we apply HyperEF (Algorithm 2) to the hypergraph $H=(V,E,w)$ to compute a coarsened hypergraph $H^{\prime}=(V^{\prime},E^{\prime},w^{\prime})$ and identify isolated (unclustered) nodes, denoted by $\mathcal{C}$. These isolated nodes are simply those that remain unclustered after HyperEF’s multilevel coarsening step. Leveraging the introduced flow-based clustering in [54], HyperSF constructs a sub-hypergraph $H^{\prime}_{L}$ by iteratively expanding the hypergraph around the seed node set $\mathcal{C}$ and then repeatedly solve the hypergraph cut problem to minimize HLC until no significant changes in local conductance are observed. Define $E^{\prime}(S)=\cup_{v^{\prime}\in V^{\prime},v^{\prime}\in\{e^{\prime}\}_{e^{\prime}\in E^{\prime}}}E^{\prime}(v)$ for any set $S\subseteq V^{\prime}$ and let $H^{\prime}_{\varsigma}=(V^{\prime}\cup\{s,t\},E^{\prime}\cup E^{\prime st})$, where $E^{\prime st}$ denotes the terminal edge set. HyperSF aims to construct a sub-hypergraph $H^{\prime}_{L}$ to replace $H^{\prime}_{\varsigma}$ that minimizes HLC by repeatedly solving a local version of (20). To this end, HyperSF sets up an oracle to discover a set of best neighborhood vertices for a given vertex $v^{\prime}$:
@@ III-C Algorithm Complexity for Spectral Coarsening
III-D Hypergraph Partitioning with Spectral Coarsening and Flow-Based Clustering
Building on previous hypergraph partitioning methods, we replace the coarsening algorithm with our proposed resistance-based approach and incorporate flow-based community detection to further enhance partitioning quality.
III-D1 Multilevel Spectral Coarsening with HyperEF
While multilevel hypergraph partitioning frameworks like KaHyPar use hyperedge size metrics in their coarsening phase to determine which vertex pairs should be contracted, we introduce a resistance-based rating function that preserves structural properties when generating a hierarchy of smaller hypergraphs. For a given vertex pair $(p,q)$, we define the rating function as:
$r(p,q)=\sum_{\begin{subarray}{c}e\in E\\ \{p,q\}\subseteq e\end{subarray}}\frac{w(e)}{R_{e}-1}$ (22)
where $w(e)$ represents the weight of hyperedge $e$, and $R_{e}$ denotes its effective resistance. The effective resistance of each hyperedge, $R_{e}$ for $e\in E$, is computed using Eq. (17) in nearly-linear time. The effective resistance information is further propagated throughout the multilevel coarsening scheme by using a node weight propagation technique, as described in Eq. (18). This approach enables us to obtain a bound on the maximum distance between any pair of nodes within a hyperedge, thereby reducing the hypergraph size while maintaining the key spectral properties of the original hypergraph.
III-D2 Flow-Based clustering with HyperSF
To further improve the quality of partitioning, we incorporate flow-based community detection using HyperSF as described in Algorithm 3, as a preprocessing step. HyperSF creates high-quality clusters by analyzing the hypergraph flow structure.
The key insight of our approach is to first identify communities using HyperSF and then apply the multilevel spectral coarsening algorithm to each community separately. This community-aware coarsening strategy produces more balanced partitions that better preserve the hypergraph’s structural properties.
HyperSF sorts clusters by their resistance diameter and prioritizes those with the lowest values. It then applies flow-based techniques to merge nearby clusters with low resistance diameters within the same neighborhood. The resulting clusters exhibit low hypergraph local conductance (HLC), which helps preserve the spectral properties of the original hypergraph during the coarsening process.
III-D3 SHyPar Algorithm Flow
In Algorithm 4, we present the detailed procedure of SHyPar, which leverages resistance-based spectral coarsening and flow-based clustering techniques for hypergraph partitioning.
The input is a weighted hypergraph $H=(V,E,w)$, and the output is a partitioned hypergraph where each node is assigned a partition index. Line 1 represents the one-time computation of effective resistance using the HyperEF algorithm, which is employed during the coarsening phase. Line 2 performs a one-time computation of flow-based clusters via the HyperSF algorithm, aiding in community detection. Lines 3–5 outline the multilevel hypergraph partitioning pipeline, including coarsening, initial cut computation, and iterative refinement. Line 6 produces the final partitioning solution of the hypergraph.
Algorithm 4 SHyPar Algorithm
Input: Hypergraph $H=(V,E,w)$
Output: Partitioned Hypergraph;
1:  Setup: Hyperedge Effective resistance computation using HyperEF algorithm 1
2:  Setup: Flow-based community detection using HyperSF algorithm 3
3:  Multilevel spectral coarsening using Eq.22
4:  Initial partitioning based on KaHyPar algorithm
5:  Solution refinement based on KaHyPar algorithm
6:  Return the partitioned hypergraph
@@ IV Experimental Validation
There are many applications related to hypergraph partitioning. This paper focuses on comprehensively evaluating the performance of the proposed hypergraph partitioning framework for increasingly important applications related to integrated circuits computer-aided design. Both the solution quality and runtime efficiency will be carefully assessed by testing them on a wide range of public-domain data sets. The implementation of the proposed algorithm, along with the code for reproducing the experimental results, is publicly available at https://github.com/Feng-Research/SHyPar.
@@ IV-B Hypergraph Partitioning with Spectral Coarsening
We compared SHyPar with leading hypergraph partitioners hMETIS [2], SpecPart [43], KaHyPar [45], and MedPart [56] using two sets of publicly available benchmarks: the ISPD98 VLSI Circuit Benchmark Suite [55] and the Titan23 Suite [57]. The details of these benchmarks are outlined in Table III and Table IV. All tests were conducted on a server equipped with Intel(R) Xeon(R) Gold 6244 processors with 1546GB of memory.
@@ IV-B2 SHyPar Performance on ISPD98 Benchmarks
TABLE III: Statistics of ISPD98 VLSI circuit benchmark suite and cut sizes obtained by different approaches. The best results among all methods are highlighted in red.
@@ IV-B3 SHyPar Performance on Titan23 Benchmarks
Building on the findings from the ISPD98 benchmarks, our research extended to the Titan23 benchmarks, notable for their high-degree hyperedges. Table IV details the results on the Titan23 benchmarks, where SHyPar dramatically outperforms hMETIS, achieving a 12% improvement for $\epsilon$ = $2\%$ and an impressive 22.5% for $\epsilon$ = $20\%$. Also SHyPar outperforms SpecPart, achieving a 3% improvement for $\epsilon$ = $2\%$ and 1% for $\epsilon$ = $20\%$. In cases such as sparcT2_core, SHyPar even exceeds the best published results by up to 15%; these notable achievements are highlighted by underlining.
TABLE IV: Statistics of the Titan23 benchmark suite and cut sizes obtained by different approaches. The best results among all methods are highlighted in red.
@@
[…]
IV-B4 SHyPar Runtime Comparison with KaHypar
Table V presents a runtime comparison between SHyPar and KaHyPar on the Titan23 benchmark for $\epsilon$ = $2\%$. The setup time corresponds to the one-time computation of the effective resistance (algorithm 1) and the flow-based clustering method (algorithm 3). The SHyPar/KaHyPar runtime ratio compares the runtime of SHyPar with the KaHyPar method, excluding the one-time setup time of effective resistance computation and flow-based clustering in SHyPar. To ensure a fair comparison, the runtime of Louvain clustering is also excluded from KaHyPar. The results demonstrate that our method achieves up to $8\%$ improvement in runtime over KaHyPar.
TABLE V: SHyPar runtime comparison with KaHyPar
Benchmark
KaHyPar
Setup Time
$\frac{\text{SHyPar}}{\text{KaHyPar}}$
sparcT1_core 52.99 10.19 0.94
neuron 41.06 10.71 0.93
stereo_vision 3.19 10.63 0.99
des90 49.90 13.29 0.94
SLAM_spheric 10.12 12.37 0.99
cholesky_mc 7.18 12.15 0.99
segmemtation 8.79 13.92 0.99
bitonic_mesh 43.52 18.04 0.93
dart 47.19 16.80 0.94
openCV 47.23 18.07 0.94
stap_qrd 22.52 19.28 0.93
minres 23.65 20.25 0.93
cholesky_bdti 35.40 20.86 0.92
denoise 22.66 23.57 0.93
sparcT2_core 49.21 22.05 0.94
gsm_switch 139.76 32.38 0.98
mes_noc 74.12 36.82 0.96
LU230 1071.02 38.26 1.00
LU_Network 78.99 43.66 0.96
sparcT1_chip2 167.14 50.63 0.98
directrf 80.89 68.07 0.96
bitcoin_miner 471.46 71.27 0.99
IV-B5 Ablation Study
To evaluate the impact of spectral coarsening and flow-based community detection, we consider a new configuration that employs only spectral coarsening within KaHyPar. Table VI presents the cut size results for some benchmarks for $\epsilon$ = $2\%$. While replacing KaHyPar’s default coarsening with spectral coarsening (SC) already outperforms the baseline, adding flow-based community detection (Flow-Based CD) on top yields even better solutions.
TABLE VI: Cut size improvement with Spectral Coarsening
Benchmark KaHyPar KaHyPar with SC
KaHyPar with
SC and Flow-Based CD
dart 924 803 784
OpenCV 560 549 499
sparcT2_core 1186 1184 1183
gsm_switch 1759 1732 1621
LU230 4012 3610 3602
IBM05 1712 1709 1707
IBM07 894 892 882
IBM08 1157 1146 1140
IBM10 1318 1272 1254
IBM18 1915 1551 1521